The potential equals 8.16 V at the midpoint between two point charges that are 1.03 m apart. One of the charges is 1.03 10-9 C. Find the value of the other charge.
Answer #1

V = V1 + V2

8.16 = (k q1/R) + (k q2/R)

q1 + q2 = 8.16 x R/k

q2 = (8.16 R/k) - q1 = (8.16 x 0.515/ 9 x 10^9) - (1.03 x 10^-9)

Charge, q2 = - 5.63 x 10^-10 C

Comment in case any doubt please rate my answer.....